150g of a gas at 10.0°C and 1.0 atm pressure occupies 913L volume. The volume of the gas increases to 988L and its temperature increases to 41.5°C while the pressure remains constant. Calculate ΔE, q and w for the gas. (Specific heat of the gas = 5.2J/g°C and 1L.atm = 101.3J).
1
Expert's answer
2020-06-08T15:22:47-0400

"W=\int PdV=P\triangle V (\because P" is constant")" "=1\times (988-913)\times 101.3 \ J=" "7597\times 10^{-3} KJ=7.6KJ"

"Q=" "mC_P\triangle T=150\times 5.2J/g\degree C\times(41.5-10)=" "24.57\ KJ\approx24.6\ KJ"

"\triangle E=Q+W" (As work is done by the gas)"=7.6+24.6=32.2\ KJ"